lamamalona:

Which equation can pair with x – y = –2 to create a consistent and dependent system?

QuestionCoveBot:

Are there more equations?

lamamalona:

6x + 2y = 15 –3x + 3y = 6 –8x – 3y = 2 4x – 4y = 6

lamamalona:

what do i do with that ??

Mercury:

a consistent and dependent solution consists of two equivalent equations, and results in a perfectly overlapping graph also recall that you can multiply an entire equation by any constant, and both sides will still be equal, resulting in two equations that have the same graph as a major hint, you can multiply x – y = –2 by the constant (-3) to get another equivalent equation that's in the answer choices